Number of integers in the range of f(x),
f(x) = 5/(5x^2+10x=8), x belongs to R

Grade:12

1 Answers

Riddhish Bhalodia
7 years ago
I believe the denominator would be5x^2 +10x +8
then then minimum value that the denominator can take is at x = -b/2a = -1 (here)

at x= -1 the denominatr becomes = 5-10+8 = 3
so the range of the denominator is[3,\infty)
thus the range of f(x) is(0,5/3]

Thus numer of integers is = 1